Swifty Company had $366600 of current assets and $141000 of current liabilities before borrowing $40400 from the bank with a 3-month note payable. What effect did the borrowing transaction have on Swifty Company's current ratio?
Answer #1

Present current ratio = Current assets/Current liabilities = 366600/141000 = 2.60

Current ratio after borrowing = (366600+40400)/(141000+40400) = 2.24

Current ratio is decline after borrowing